OpenStudy (anonymous):

multiply (x^9+ x^8)(x^9 + x^8 -x^7)

OpenStudy (blockcolder):

For a brief moment, let u=x^9+x^8. u(u-x^7)=u^2-u(x^7)=x^18+2x^17+x^16-x^16-x^15=x^18+2x^17-x^15